OpenStudy (anonymous):

Give an equation of the line through the point (4, 1) and perpendicular to the line x + 5y = 1.

OpenStudy (jdoe0001):

hmmm what do you think is the slope of => x + 5y = 1. ?

OpenStudy (anonymous):

-1/5?

OpenStudy (jdoe0001):

yeap is -1/5 so, the slope of a line perpendicular to that one, will be the NEGATIVE RECIPROCAL of that, that is \(\bf -\cfrac{1}{{\color{blue}{ 5}}}\qquad negative\to +\cfrac{1}{{\color{blue}{ 5}}}\qquad reciprocal\to +\cfrac{{\color{blue}{ 5}}}{1}\to 5\) so you're really asked on how to find the equation of a line with slope of 5 and that passes through (4,1) that is \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 1}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 5 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

y-1=5(x-(-1)

OpenStudy (anonymous):

y=5x+1?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 1}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 5 \\ \quad \\ y-{\color{blue}{ 1}}={\color{green}{ 5}}(x-{\color{red}{4}})\implies y-1=5x-20\implies y=5x-20+1\)

OpenStudy (anonymous):

So that would be y=5x-19?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks! but should it be in standard form?

OpenStudy (jdoe0001):

well... it doesn't say above...so either will do I'd think, that's just the slope-intercept one, but you can put it in standard

OpenStudy (anonymous):

Thanks! (-:

OpenStudy (jdoe0001):

yw